# Jacobian-Enhanced Neural Networks

### v1

@@ Acknowledgement
The supporting code for this work used the exercises by Prof. Andrew Ng [14] as a starting point. It then built upon it to include additional features but, most of all, it fundamentally changed the formulation to include jacobian-enhancement and made sure all arrays were updated in place (data is never copied). The author would therefore like to thank Andrew Ng for offering the fundamentals of deep learning on Coursera, which took a complicated subject and explained it in simple terms that even an aerospace engineer could understand.
@@ VIII Key Equations in Vectorized Form
Key equations from the preceding material will now be re-written in a vectorized form (i.e. processing all examples simultaneously without loops) suitable for programming, as a helpful companion to source code. In the notation to follow, the symbol implies element-wise array multiplication.

### v2

@@ Acknowledgement
The supporting code for this work used the exercises by Professor Andrew Ng [23] as a starting point. It then built upon it to include additional features but, most of all, it fundamentally changed the formulation to include jacobian-enhancement and made sure all arrays were updated in place (data is never copied). The author would therefore like to thank Professor Andrew Ng for offering the fundamentals of deep learning on Coursera, which masterfully explained what can be considered a challenging and complicated subject in an easy to understand form.
@@ VIII Key Equations in Vectorized Form
This section provides vectorized equations to facilitate re-implementation or as a companion guide to the JENN library.